jagr2713 (jagr2713):

\frac{ 6x-7 }{ 4 }+\frac{ 3x-5 }{ 7 }=\frac{ 5x+78 }{ 28 }

jagr2713 (jagr2713):

\[\frac{ 6x-7 }{ 4 }+\frac{ 3x-5 }{ 7 }=\frac{ 5x+78 }{ 28 }\]

jagr2713 (jagr2713):

i got x=3

jagr2713 (jagr2713):

@sleepyjess

TheSmartOne (thesmartone):

to make it pop up in the top just add `\(` and `\)` that in the top :P

OpenStudy (sleepyjess):

Correct

jagr2713 (jagr2713):

oh ok and thanks